OpenStudy (anonymous):

the solution to the system y=2x-3 is: 5x=2y=30

OpenStudy (imstuck):

Do you have a question here?

OpenStudy (anonymous):

I think he's asking what is the solution to that system of equation is that right?

OpenStudy (anonymous):

yes ,that is right.

OpenStudy (anonymous):

There are many methods to solve this but you can start by using substitution. y=2x-3 Is the second equation 5x=2y-30? That = sign doesn't belong there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The first equation already gives you what y= to. It equals to 2x-3. You can just plug y=2x-3 into 5x=2y-30. Your equation would then be 5x=(2(2x-3))-30 Now you should solve for x after simplifying 5x= 4x-6-30 5x=4x-36 Subtract 4x on both sides x=-36 Plug x=-36 back into one of the equations above. For example, y=2x-3 2(-36)-3 = -75 That means your answer is (-36, -75) Plug those values into your equation and they should both work.